Prove that every first countable space is second countable


1
Expert's answer
2021-07-18T18:19:43-0400

Let "(X,d)" be a metric space and let "x\in X". Consider the neighborhood basis

"\mathbb{B}_x=\{B_r(x) | r>0,r\in \mathbb{Q}\}".

Clearly, this basis is countable. Hence by definition, every metric space is first countable.


Now let "(\mathbb{R},d)" be a metric space where "d" is the discrete metric

"d(x,y) =
 \begin{cases}
 0 & \text{if $x=y$} \\
 1 & \text{if $x\neq y$} 
 \end{cases}"

Clearly, a base for the topology induced by this metric must contain all singleton subsets of "\mathbb{R}". So it is uncountable (since "\mathbb{R}" is uncountable). Hence "(\mathbb{R},d)" is not second countable.